An electric heater is to be made from a

nichrome wire. Nichrome has a Resistivity of 1.0*10^-6 ohm'sm

at the operating temperature of

the heater. The power of the heater is 60W

when the potential difference across its

terminals is 12V. Calculate the length of the

wire of diameter 0.30mn required for the

heater.


1
Expert's answer
2022-03-14T13:06:17-0400

Answer

Length of wire is given by

"L=\frac{V^2A}{\rho P}\\=\frac{12*12*3.14*0.15*10^{-3}}{1.0*10^{-6}*60}\\=6782.4m"